prettygirlLA:

Evaluate 4z-5y y=3 z=8

supie:

So what you have to do is plug in the values of the variables So right now it is \(4z-5y\) which is 4 time z minus 5 times y So when we plug in the numbers we get \(4 \times 8-5 \times 3\)

prettygirlLA:

ummmm yea no

supie:

? You don't understand...?

prettygirlLA:

you dont times it bye 8 or 3 it would be 4x4x4x4x4x4x4x4 and 5x5x5

prettygirlLA:

by*

supie:

So it is supposed to be \(4^z-5^y\)?

prettygirlLA:

yes

supie:

...well it's not my fault that you didn't include that...

supie:

the way you write that is use a `^`

supie:

So all you have to do is plug in the values of y and z

prettygirlLA:

hold on no the letters at the bottom

supie:

....what?

prettygirlLA:

dont get math i thought it was the same thing if thy were on top or bottom

supie:

Ok so all you have to do is plug in the values of y and z \(4^z-5^y\longrightarrow 4^8-5^3\) So then all you have to do is simplify \(4^8-5^3\)

supie:

Ok so the first thing we do is solve the first exponent \(\color{red}{4^8}-5^3\) \(4\times4\times4\times4\times4\times4\times4\times4\)

CripQUEZZ:

I thought it would jus be (4 x 3) - (5 x 8)

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @CripQUEZZ I thought it would jus be (4 x 3) - (5 x 8) \(\color{#0cbb34}{\text{End of Quote}}\) Yes I thought that also but she forgot to include that `z` and `y` are exponents so its not \(4z-5y\) it's \(4^z-5^y\)

CripQUEZZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @CripQUEZZ I thought it would jus be (4 x 3) - (5 x 8) \(\color{#0cbb34}{\text{End of Quote}}\) I mean (4 x 8) - (5 x 3)

supie:

They are exponents,

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie Ok so the first thing we do is solve the first exponent \(\color{red}{4^8}-5^3\) \(4\times4\times4\times4\times4\times4\times4\times4\) \(\color{#0cbb34}{\text{End of Quote}}\) So then we do the second one, \(5\times5\times5=?\)

CripQUEZZ:

Ok so (4^8) - (5^3) = So then they 4x4x4x4x4x4x4x4 - 5x5x5 = So then they simplify to 65,536 - 125 = And they will get 65,411

CripQUEZZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @CripQUEZZ Ok so (4^8) - (5^3) = So then they 4x4x4x4x4x4x4x4 - 5x5x5 = So then they simplify to 65,536 - 125 = And they will get 65,411 \(\color{#0cbb34}{\text{End of Quote}}\) I think that’s it

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie \(\color{#0cbb34}{\text{Originally Posted by}}\) @supie Ok so the first thing we do is solve the first exponent \(\color{red}{4^8}-5^3\) \(4\times4\times4\times4\times4\times4\times4\times4\) \(\color{#0cbb34}{\text{End of Quote}}\) So then we do the second one, \(5\times5\times5=?\) \(\color{#0cbb34}{\text{End of Quote}}\) So basically \(4^z-5^y\longrightarrow 4^8-5^3\longrightarrow 4\times4\times4\times4\times4\times4\times4\times4-5\times5\times5\longrightarrow~?\) So find what \(4\times4\times4\times4\times4\times4\times4\times4\) and \(\times4-5\times5\times5\) equals, then subtract the product of both of those.

CripQUEZZ:

Wait so I am right?

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @CripQUEZZ Wait so I am right? \(\color{#0cbb34}{\text{End of Quote}}\) Yes you're right nicely explained.

CripQUEZZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie \(\color{#0cbb34}{\text{Originally Posted by}}\) @CripQUEZZ Wait so I am right? \(\color{#0cbb34}{\text{End of Quote}}\) Yes you're right nicely explained. \(\color{#0cbb34}{\text{End of Quote}}\) Ok thank you

supie:

np